OpenStudy (anonymous):

please help me solve!...differentiate y=(3x+1)^2

OpenStudy (lgbasallote):

use chain rule

OpenStudy (ganpat):

or simplify the equation and than solve individually...

OpenStudy (lgbasallote):

here's a hint (i'll give a demo) derivative of \[(3x+2)^3 = 3(3x+2)^2 \times 3 \implies 9(3x+2)^2\] does that help?

OpenStudy (anonymous):

ya that does help...so either chain rule of simplifying will work?

OpenStudy (anonymous):

i got 6(3x+1) is that right??

OpenStudy (ganpat):

y=(3x+1)^2 = 9x2 + 6x + 1 and differentiate now.. so y` = 18x + 6... is that ur answer ??

OpenStudy (anonymous):

ya..is it terribly wrong?

OpenStudy (ganpat):

we got the same dude... @mathdummy4 ... @lgbasallote : can u plz check the answer !!

OpenStudy (anonymous):

o you just multiplied yours out one more time...oops lol

OpenStudy (lgbasallote):

yeah both are right

OpenStudy (anonymous):

thanks!!